# Optimized Current Density Reconstruction from Widefield Quantum Diamond Magnetic Field Maps

### v1

@@ Abstract
In this work, we conduct an in-depth comparative analysis of the Fourier-based and Bayesian reconstruction methods for both simulated and experimental magnetic maps obtained from a quantum diamond microscope where the sample is assumed to be placed at a standoff distance $z$ from the source, as schematically outlined in Fig. 1(a). The reconstruction pipeline is further outlined in Fig. 1(b). We perform a thorough examination of the performance of reconstruction algorithms through extensive numerical simulations within various sensor regimes characterized by noise levels and standoff distances in Sec. I. Following this, in Sec. II we build on our theoretical analysis by exploring practical approaches for determining the regularizer strength in Bayesian algorithms, and applying them to experimental data collected in three different noise and standoff scenarios. Through this, we thereby introduce the concept of Bayesian advantage in the realm of current density reconstructions for NV widefield experiments.
I Simulation results and discussions
As a test structure, we use a microcoil wire with multiple sharp bends to effectively highlight nuances of reconstruction algorithms. A comprehensive outline of the simulation method, which includes the process for computing magnetic fields from current densities, as well as a detailed explanation of the reconstruction methodology is provided in the Supplementary Material. The ground truth current density is depicted in Fig. 1(c), where we define the $j_{x}$ and $j_{y}$ components of the spiral wire piecewise in a total image size of $50\times 50\mu m^{2}$. The scale bar in Fig. 1 is $10\mu m$. We assume a constant current density of $100A/m$ across the wire. The resulting magnetic field, computed at a standoff of $z=1\mu m$, using a Fourier-space application of the Biot-Savart kernel $\mathcal{M}$, is shown in Fig. 1(d).
@@ I Simulation results and discussions
The reconstructed $j$-maps, upon application of both the Fourier and Bayesian methods on the magnetic images in Fig. 1(d) are shown in Fig. 1(e). The regularizer strength $\lambda=\lambda^{*}$ is carefully selected in the Bayesian method, following a systematic and experimentally-informed procedure that we outline in the Supplementary Material. In the rest of the simulation results, we perform the comparative analysis of the Bayesian method in a non-adaptive setting without performing regularizer selection, to outline worst case behavior. The following three sections outline the analysis of the three reconstruction algorithms on the simulated data in various noise and standoff regimes.
I.1 The Effect of Standoff
@@
The distance between the sample and the NV sensor, or the “standoff” $(z)$, is a key parameter that strongly affects the measured magnetic field maps and consequently the underlying current distribution. This can be seen from the exponential dependence on standoff distance in the Biot-Savart kernel, $\mathcal{M}$ (See Supplementary Material). This underscores the need for a thorough analysis of how the performance of the algorithms is influenced by the standoff distance.
@@
We plot two error-estimation metrics, the two-normed distance ($D_{L2}$) and structural similarity index (SSIM), as a function of standoff $z$ in Fig. 2(a) and (b) respectively, without any added noise ($\eta=0$). The two-normed distance measures the Euclidean distance between two images, and the SSIM measures a ‘score’ of visual similarity (See Supplementary Material). The Bayesian reconstructions require selecting the strength of regularizer term, which maintains smoothness of the reconstructed solution. Here, we do not perform systematic regularizer selection and rather choose three fiducial strengths $\lambda\in\{0.01,0.1,1\}$ in a non-adaptive fashion. The Fourier methods on the other hand require choosing a cut-off frequency to band-limit the spectrum and perform inversions (see Supplementary Material). We perform adaptive Fourier reconstructions ($\mathcal{F}_{XY}$, $\mathcal{F}_{Z}$) with the cutoff frequency of the algorithms changing with every $2\mu m$ increase in standoff, as elaborated in the Supplementary section. Two points representing the low standoff $z_{1}\lesssim 1\mu m$ and high standoff $z_{2}\gtrsim 5\mu m$ are marked in the error plots with dotted lines. First, we observe that in both $D_{2L}$ and SSIM, the Fourier-XY method ($\mathcal{F}_{XY}$) performs better for the low-standoff regime of $z\lesssim 1\mu m$. As the standoff is increased, there exists a critical point beyond which the Bayesian method ($\mathcal{B}$) method starts to outperform $\mathcal{F}_{XY}$. This can be explained by noting that the $\mathcal{F}_{XY}$ method is exact at low stand-off, eliminating aberrations due to its ability to access both in-plane components, however, the spiral wire contains a non-trivial frequency spectrum owing to multiple sharp corners, and thereby requires a careful selection of a band-limiting frequency for the Fourier algorithms at higher standoff. Additionally, without such a selection of the cutoff frequency in the Fourier algorithms at higher standoffs, high-frequency artifacts dominate the Fourier reconstruction. This leads to a sharp unbounded increase in the $D_{2L}$ metric at $z\approx 8\mu m$. In contrast, the Bayesian reconstruction maintains a near-constant reconstruction quality for standoff distances of $z\lesssim 5\mu m$.
@@
I.2 Robustness to Noise
We now study the effect of noise on reconstruction. We employ two noise models—Gaussian noise, which distributes across the entire image uniformly, and scaled noise, which is more prominent at pixels with higher absolute values. Both the models add zero-mean Gaussian noise to each pixel, with the former having a pixel independent variance and the latter having a variance that scales with the absolute value of the pixel. The amount of noise in both models is controlled by a single parameter $\eta$, and we defer the details to the Supplementary Material. We study the SSIM distance between the reconstructed and ground truth current densities as a function of noise strength $\eta$. Results are shown here for Gaussian noise, with scaled noise exhibiting similar patterns—differences between the two noise models are shown in the next subsection. Figure 3 shows the results for (a-b) the low-standoff regime of $z=0.1\mu m$ and (c-d) the high standoff regime of $z=5\mu m$. For Bayesian reconstructions, $\lambda\in\{0.1,1.0\}$ is chosen. In the scenario with a low standoff, the $\mathcal{F}_{XY}$ method emerges as the most effective, displaying the highest SSIM score for the entire noise range $0\leq\eta_{g}\leq 2$. The SSIM also reveals two distinct patterns at low standoff — a rapid decline in algorithm performance with increasing noise, followed by a shift in slope after reaching a critical noise strength. This phenomenon can be linked to a critical signal-to-noise (SNR) ratio, where the noise levels equate to the signal strength, resulting in two performance regimes. The $\mathcal{F}_{Z}$ image exhibits a low SNR ratio within the wire, leading to diminished SSIM as noise levels escalate. In contrast, the $\mathcal{B}(\lambda)$ reconstructions exhibit high SNR, as well as the characteristic offset associated with the lack of regularizer selection.
@@ I.3 Phase Diagram
The outcomes for $d(z,\eta)$ are presented in Fig. 4 for (a) Gaussian and (b) scaled noise, with sub-panels illustrating the model selection for the Bayesian approach using $\lambda\in{0.1,1}$ in accordance with the earlier discussions. The dashed line signifies a distinctive phase transition from Fourier to Bayesian advantage, for which $d(z,\eta)$ serves as an effective order parameter. In the case of Gaussian noise in Fig. 4(a), results indicate that at very low standoffs, there is no clear preference for either of the two methods, as was seen in Fig. 3(a). Additionally, for $\lambda=1$ in the Bayesian scenario, the Fourier advantage region diminishes, suggesting that adequate regularization in Bayesian optimization generally leads to improved reconstruction quality without the need for regularizer tuning. In the case of scaled noise in Fig. 4(b), similar to a ‘phase transition,’ the results show a notable contrast with a pronounced Fourier advantage at low standoffs, as evidenced by the image contrast. Moreover, the area of Fourier advantage is smaller in the Gaussian noise case (Fig. 4(a)) as compared to the case of scaled noise (Fig. 4(b)). This is because of the fact that Gaussian noise has a flat and non-decaying spectrum for any $\eta_{g}>0$, playing a more detrimental effect in Fourier inversions.
@@
II Reconstruction from Experimental Magnetic Maps
We now apply the reconstruction algorithms to three chosen experimental scenarios to illustrate the points described via simulations in the previous section. The regularizer selection—the search of the optimal trade-off $\lambda^{*}$—for Bayesian reconstructions on the experimental data is performed using the $l$-curve method and via an analysis of the total integrated currents. Details of the selection method are outlined for the simulated micro-coil wire in the Supplementary Material.
II.1 Micro-spiral wire
@@ IV Acknowledgments
First, the simulation methodology is introduced, followed by a discussion about the reconstruction algorithms and the error-estimation metrics employed in the main text. Then, a comprehensive explanation of the regularizer selection in the Bayesian algorithm for the spiral wire case is presented. Lastly, details of the experiments are discussed.
IV.1 Simulated Magnetic Fields
@@
The magnetic field simulations are performed by (i) specifying the (two-dimensional) wire structure and magnitude of the current density map, (ii) applying the Biot-Savart law, and subsequently (iii) adding suitable noise. Aiming to analyze a structure with non-trivial features, we work with a micro-coil wire with multiple sharp turns aiming to highlight issues with the reconstruction algorithms. The current density across this wire is held constant at $100A/m$, and the $x$ and $y-$components of the current density are specified piecewise in each arm of the wire. We compute the fields resulting from this wire at a standoff of $z~{}\mu m$ using a Fourier-space application of the Biot-Savart kernel $\mathcal{M}$. Now, as any experimental data contains noise, we add noise to these simulated images to (i) emulate experimental conditions and (ii) evaluate the noise robustness of the algorithms. The noise is added in two regimes – (i) Gaussian noise and (ii) scaled noise. The amount of noise under both the regimes is controlled by a single parameter $\eta$, as follows: Given a matrix $A\equiv A(i,j)\in\mathbb{R}^{n\times m}$ we add noise as follows:
@@ IV.2 Reconstruction Methodologies
The results upon application of the three reconstruction algorithms on the spiral wire for the no-noise case of standoff $1\mu m$ are shown in Fig. 1 of the main text.
IV.3 Error estimation
@@
IV.4 Regularizer Selection in Bayesian Reconstructions
Supplementary Figure 1: Procedure for selecting the regularizer strength for a constant Gaussian noise with strength $\eta_{g}=0.5$. (a) Results for the small standoff regime of $z=0.1\mu m$. (b) Results for the high standoff of $z=5\mu m$. Subpanels of (a-b): (i) Noise magnetic field. (ii) $l$-curve with three regularizer strengths marked for inspection. (iii) The residual distance between the clean and reconstructed magnetic field (left axis) and the SSIM distance between the true and reconstructed densities (right axis). (iv) Reconstructions for three cases of regularizer strengths marked in the $l$-curve.
@@ IV.4 Regularizer Selection in Bayesian Reconstructions
and outputs $\lambda=\lambda^{*}$ if the noise in the residual matches with the actual noise of the sensing, i.e., $\text{std}(r_{\lambda})=\sigma$. The challenge with such approaches lies in the complexity of determining the noise level, denoted as $\sigma$, in a widefield NV experiment. This complexity necessitates a method that is not reliant on $\sigma$. We employ the $l$-curve approach [33] in this work for selecting the regularizer strength, establishing its significance as a experimentally-relevant method. This method involves plotting a parameterized curve representing the norm of the residual $||\mathcal{M}(\hat{j}_{\lambda})-B||_{2}$ against the norm of the reconstruction $||\hat{j}||_{2}$. This visualization explores the trade-off between fitting the data and regularizing the solution, revealing a distinct ‘kink’ in the curve that corresponds to the correct regularizer strength. Although this method is originally analytical for Tikhonov regularization, where the norm of the reconstruction is regulated, we demonstrate its effectiveness for total variation regularization as well. To support this, we include a plot of the SSIM$(j,\hat{j})$.
@@
The results are shown in Supp. Fig.1, for (a) low standoff of $z=0.1\mu m$ and (b) high standoff of $z=1\mu m$. The panels show (i) the noisy magnetic field, (ii) the $l-$curve, (iii) a plot of the residual norm as well as the SSIM score, and (iv) some reconstructed densities corresponding to $\lambda$-points marked in the $l-$curve. The noise level is kept constant at $\eta_{g}=0.5$ for both the standoff regimes.
@@
The $l$-curve in Supp. Fig.1(a) shows a turn at $\lambda=2.04$, with two additional points marked at $\lambda\in\{1.22,2.53\}$. The value $\lambda=2.53$ was chosen in line with the plot of SSIM as a function of $\lambda$ as shown in Supp. Fig.1(a)(iii), and corresponds to the maximum similarity distance between the reconstructed and true densities. We show the reconstructed densities at these three strengths in Supp. Fig.1(a)(iv), and remark that $\lambda=1.22$ shows noise artifacts inside and outside the wire and is thus before the turn in the $l-$curve. The $\lambda=2.04$ shows no artifacts, but has a higher offset. Instead, the maxima of the SSIM curve at $\lambda=2.53$ shows a lower offset with no artifacts outside the wire. This demonstrates that the $l-$curve is not exact for TV regularization and should instead be used in supplement with SSIM to choose the regularizer strength.
@@
The outcomes for a high standoff are presented in Supp. Fig.1(b). The $l$-curve exhibits a distinct inflection point at $\lambda=0.2$, with two supplementary points $\lambda\in\{0.06,0.41\}$ marked for reference. However, the SSIM curve displays a noticeable contrast compared to the low standoff scenario. Instead of a peak, there is a plateau, and a reference local maximum is identified on this plateau at $\lambda=0.41$. Additionally, as anticipated, the SSIM distance is lower than in the low standoff case. This indicates that the SSIM curve alone is insufficient for regularizer selection, as all models with $\lambda\in[0.2,0.5]$ yield the same SSIM, rendering it challenging to make a selection. This issue is further elucidated by the reconstructions in Supp. Fig.1(b)(iv). The image corresponding to $\lambda=0.2$, identified by the kink in the $l$-curve, demonstrates the optimal balance between resolving wire features and suppressing noise. In contrast, the $\lambda=0.06$ image exhibits excessive noise, while the $\lambda=0.41$ plot shows significant broadening. Consequently, employing the $l$-curve in conjunction with an additional distance metric like SSIM offers a reliable method for selecting the regularizer strength in Bayesian optimization.

### v2

@@ I Introduction
In this work, we conduct an in-depth comparative analysis of the Fourier-based and Bayesian reconstruction methods for both simulated and experimental magnetic maps obtained from a quantum diamond microscope where the sample is assumed to be placed at a standoff distance $z$ from the source, as schematically outlined in Fig. 1(a). The reconstruction pipeline is further outlined in Fig. 1(b). We perform a thorough examination of the performance of reconstruction algorithms through extensive numerical simulations within various sensor regimes characterized by noise levels and standoff distances in Sec. II. Following this, in Sec. III we build on our theoretical analysis by exploring practical approaches for determining the regularizer strength in Bayesian algorithms, and applying them to experimental data collected in three different noise and standoff scenarios. Through this, we thereby introduce the concept of Bayesian advantage in the realm of current density reconstructions for NV widefield experiments.
II Simulation results and discussions
As a test structure, we use a microcoil wire—carrying a purely two-dimensional (2D) current to model a quasi 2D flow—with multiple sharp bends to effectively highlight nuances of reconstruction algorithms. A comprehensive outline of the simulation method, which includes the process for computing magnetic fields from current densities, as well as a detailed explanation of the reconstruction methodology is provided in the Supplementary Material [32]. The assumed true current density, which by definition forms the ground truth is depicted in Fig. 1(c), where we define the $j_{x}$ and $j_{y}$ components of the spiral wire piecewise in a total image size of $50\times 50\mu m^{2}$. The scale bar in Fig. 1 is $10\mu m$. We assume a constant current density of $100A/m$ across the wire. The resulting magnetic field, computed at a standoff of $z=1\mu m$, using a Fourier-space application of the Biot-Savart kernel $\mathcal{M}$, is shown in Fig. 1(d).
@@ II Simulation results and discussions
The reconstructed $j$-maps, upon application of both the Fourier and Bayesian methods on the magnetic images in Fig. 1(d) are shown in Fig. 1(e). The regularizer strength $\lambda=\lambda^{*}$ is carefully selected in the Bayesian method, following a systematic and experimentally-informed procedure that we outline in the Supplementary Material [32]. In the rest of the simulation results, we perform the comparative analysis of the Bayesian method in a non-adaptive setting without performing regularizer selection, to outline worst case behavior. The following three sections outline the analysis of the three reconstruction algorithms on the simulated data in various noise and standoff regimes.
II.1 The Effect of Standoff
@@
The distance between the sample and the NV sensor, or the “standoff” $(z)$, is a key parameter that strongly affects the measured magnetic field maps and consequently the underlying current distribution. This can be seen from the exponential dependence on standoff distance in the Biot-Savart kernel, $\mathcal{M}$ [32]. This underscores the need for a thorough analysis of how the performance of the algorithms is influenced by the standoff distance.
@@
We plot two error-estimation metrics, the two-normed distance ($D_{L2}$) and structural similarity index (SSIM), first introduced in [33], as a function of standoff $z$ in Fig. 2(a) and (b) respectively, without any added noise ($\eta=0$). See [34, 35] for brief reviews of the error and similarity metrics. The two-normed distance measures the Euclidean distance between two images, and the SSIM measures a ‘score’ of visual similarity [32]. The Bayesian reconstructions require selecting the strength of regularizer term, which maintains smoothness of the reconstructed solution. Here, we do not perform systematic regularizer selection and rather choose three fiducial strengths $\lambda\in\{0.01,0.1,1\}$ in a non-adaptive fashion. The Fourier methods on the other hand require choosing a cut-off frequency to band-limit the spectrum and perform inversions [32]. We perform adaptive Fourier reconstructions ($\mathcal{F}_{XY}$, $\mathcal{F}_{Z}$) with the cutoff frequency of the algorithms changing with every $2\mu m$ increase in standoff, as elaborated in [32]. Two points representing the low standoff $z_{1}\lesssim 1\mu m$ and high standoff $z_{2}\gtrsim 5\mu m$ are marked in the error plots with dotted lines. First, we observe that in both $D_{2L}$ and SSIM, the Fourier-XY method ($\mathcal{F}_{XY}$) performs better for the low-standoff regime of $z\lesssim 1\mu m$. As the standoff is increased, there exists a critical point beyond which the Bayesian method ($\mathcal{B}$) method starts to outperform $\mathcal{F}_{XY}$. This can be explained by noting that the $\mathcal{F}_{XY}$ method is exact at low stand-off, eliminating aberrations due to its ability to access both in-plane components, however, the spiral wire contains a non-trivial frequency spectrum owing to multiple sharp corners, and thereby requires a careful selection of a band-limiting frequency for the Fourier algorithms at higher standoff. Additionally, without such a selection of the cutoff frequency in the Fourier algorithms at higher standoffs, high-frequency artifacts dominate the Fourier reconstruction. This leads to a sharp unbounded increase in the $D_{2L}$ metric at $z\approx 8\mu m$. In contrast, the Bayesian reconstruction exhibits a SSIM decrease less than $10\%$ for standoff distances of $z\lesssim 5\mu m$.
@@
II.2 Robustness to Noise
We now study the effect of noise on reconstruction. We employ two noise models—Gaussian noise, which distributes across the entire image uniformly, and scaled noise, which is more prominent at pixels with higher absolute values. Both the models add zero-mean Gaussian noise to each pixel, with the former having a pixel independent variance and the latter having a variance that scales with the absolute value of the pixel. The amount of noise in both models is controlled by a single parameter $\eta$ [32]. We study the SSIM distance between the reconstructed and ground truth current densities as a function of noise strength $\eta$. Results are shown here for Gaussian noise, with scaled noise exhibiting similar patterns—differences between the two noise models are shown in the next subsection. Figure 3 shows the results for (a-b) the low-standoff regime of $z=0.1\mu m$ and (c-d) the high standoff regime of $z=5\mu m$. For Bayesian reconstructions, $\lambda\in\{0.1,1.0\}$ is chosen. The $\lambda=0.01$ case is omitted as the presence of noise warrants non-trivial regularization. In the scenario with a low standoff, the $\mathcal{F}_{XY}$ method emerges as the most effective, displaying the highest SSIM score for the entire noise range $0\leq\eta_{g}\leq 2$. The SSIM also reveals two distinct patterns at low standoff — a rapid decline in algorithm performance with increasing noise, followed by a shift in slope after reaching a critical noise strength. This phenomenon can be linked to a critical signal-to-noise (SNR) ratio, where the noise levels equate to the signal strength, resulting in two performance regimes. The $\mathcal{F}_{Z}$ image exhibits a low SNR ratio within the wire, leading to diminished SSIM as noise levels escalate. In contrast, the $\mathcal{B}(\lambda)$ reconstructions exhibit high SNR, as well as the characteristic offset associated with the lack of regularizer selection.
@@ II.3 Phase Diagram
The outcomes for $d(z,\eta)$ are presented in Fig. 4 for (a) Gaussian and (b) scaled noise, with sub-panels illustrating the model selection for the Bayesian approach using $\lambda\in\{0.1,1\}$ in accordance with the earlier discussions. The dashed line signifies a distinctive phase transition (in the spirit of similar phase diagrams in optimization dynamics, for instance see [36]) from Fourier to Bayesian advantage, for which $d(z,\eta)$ serves as an effective order parameter. In the case of Gaussian noise in Fig. 4(a), results indicate that at very low standoffs, there is no clear preference for either of the two methods, as was seen in Fig. 3(a). Additionally, for $\lambda=1$ in the Bayesian scenario, the Fourier advantage region diminishes, suggesting that adequate regularization in Bayesian optimization generally leads to improved reconstruction quality without the need for regularizer tuning. In the case of scaled noise in Fig. 4(b), similar to a ‘phase transition,’ the results show a notable contrast with a pronounced Fourier advantage at low standoffs, as evidenced by the image contrast. Moreover, the area of Fourier advantage is smaller in the Gaussian noise case (Fig. 4(a)) as compared to the case of scaled noise (Fig. 4(b)). This is because of the fact that Gaussian noise has a flat and non-decaying spectrum for any $\eta_{g}>0$, playing a more detrimental effect in Fourier inversions.
@@
III Reconstruction from Experimental Magnetic Maps
In the previous discussions, we demonstrated how inference-based protocols exhibit an advantage in current density reconstruction by studying simple uncorrelated noise. In this section, we further show that the same models indeed work on experimental data, presumably incurring non-analytical and correlated noise, and further exemplify the Bayesian advantage. We now apply the reconstruction algorithms to three chosen experimental scenarios to illustrate the points described via simulations in the previous section. The regularizer selection—the search of the optimal trade-off $\lambda^{*}$—for Bayesian reconstructions on the experimental data is performed using the $l$-curve method and via an analysis of the total integrated currents. The $l-$curve method involves searching for a bending point in a parametrized plot of the norm of the reconstructed density $||\hat{j}_{\mathcal{B}}||_{2}$ against the norm of the residual $||\mathcal{M}(\hat{j}_{\mathcal{B}})-B_{c}||_{2}$ as the $\lambda$ parameter is varied (see [37, 38]). The key idea behind selecting a near-optimal regularizer is to include correct prior information about the expected current densities in the optimization process. Mathematically, we inforce a particular functional form, known as the total-variation regularizer [39, 40] which ensures smootheness of the reconstruction by penalizing the first derivatives. The strength of this prior is then controlled by the parameter $\lambda$. A small $\lambda\approx 0$ fits to the data, which can include the noise, and will produce exactly the same results as the Fourier inversion in the case of no noise ($\eta=0$). The ingenuity of the Bayesian methods lies in tuning the $\lambda$, to be able to control the tradeoff between fitting the exact noisy data and rejecting the noise to reconstruct the correct fields. Details of the regularizer selection method are outlined with the simulated micro-coil wire as an instructive example in the Supplementary Material [32].
III.1 Micro-spiral wire
@@ V Acknowledgments
First, the simulation methodology is introduced, followed by a discussion of the reconstruction algorithms and the error-estimation metrics employed in the main text. Then, a comprehensive explanation of the regularizer selection in the Bayesian algorithm for the spiral wire case is presented. Lastly, details of the experiments are discussed.
VI Simulated Magnetic Fields
@@
The magnetic field simulations are performed by (i) specifying the (two-dimensional) wire structure and magnitude of the current density map, (ii) applying the Biot-Savart law, and subsequently (iii) adding suitable noise. Aiming to analyze a structure with non-trivial features and highlight nuances of the reconstruction algorithms, we work with a micro-coil wire with multiple sharp turns. The current density across this wire is held constant at $100A/m$, and the $x$ and $y-$components of the current density are specified piecewise in each arm of the wire. We compute the fields resulting from this wire at a standoff of $z~{}\mu m$ using a Fourier-space application of the Biot-Savart kernel $\mathcal{M}$. Now, as any experimental data contains noise, we add noise to these simulated images to (i) emulate experimental conditions and (ii) evaluate the noise robustness of the algorithms. The noise is added in two regimes – (i) Gaussian noise and (ii) scaled noise. The amount of noise under both the regimes is controlled by a single parameter $\eta$, as follows: Given a matrix $A\equiv A(i,j)\in\mathbb{R}^{n\times m}$ we add noise as follows:
@@ VII Reconstruction Methodologies
[…]
VIII Error estimation
@@
IX Regularizer Selection in Bayesian Reconstructions
Supplementary Figure 1: Procedure for selecting the regularizer strength for a constant Gaussian noise with strength $\eta_{g}=0.5$. (a) Results for the small standoff regime of $z=0.1\mu m$. (b) Results for the high standoff of $z=5\mu m$. Subpanels of (a-b): (i) Noise magnetic field. (ii) $l$-curve with three regularizer strengths marked for inspection. (iii) The residual distance between the clean and reconstructed magnetic field (left axis) and the SSIM distance between the true and reconstructed densities (right axis). (iv-vi) Reconstructions for three cases of regularizer strengths marked in the $l$-curve.
@@ IX Regularizer Selection in Bayesian Reconstructions
and outputs $\lambda=\lambda^{*}$ if the noise in the residual matches with the actual noise of the sensing, i.e., $\text{std}(r_{\lambda})=\sigma$. The challenge with such approaches lies in the complexity of determining the noise level, denoted as $\sigma$, in a widefield NV experiment. This complexity necessitates a method that is not reliant on $\sigma$. We employ the $l$-curve approach [37, 38, 39] in this work for selecting the regularizer strength, establishing its significance as a experimentally-relevant method. This method involves plotting a parameterized curve representing the norm of the residual $||\mathcal{M}(\hat{j}_{\lambda})-B||_{2}$ against the norm of the reconstruction $||\hat{j}||_{2}$. This visualization explores the trade-off between fitting the data and regularizing the solution, revealing a distinct ‘kink’ in the curve that corresponds to the correct regularizer strength. Although this method is originally analytical for Tikhonov regularization [41], where the norm of the reconstruction is regulated, we demonstrate its effectiveness for total variation regularization as well. To support this, we include a plot of the SSIM$(j,\hat{j})$.
@@
The results are shown in Supp. Fig.1, for (a) low standoff of $z=0.1\mu m$ and (b) high standoff of $z=1\mu m$. The panels show (i) the noisy magnetic field, (ii) the $l-$curve, (iii) a plot of the residual norm as well as the SSIM score, and (iv-vi) some reconstructed densities corresponding to $\lambda$-points marked in the $l-$curve. The noise level is kept constant at $\eta_{g}=0.5$ for both the standoff regimes.
@@
The $l$-curve in Supp. Fig.1(a) shows a turn at $\lambda=2.04$, with two additional points marked at $\lambda\in\{1.22,2.53\}$. The value $\lambda=2.53$ was chosen in line with the plot of SSIM as a function of $\lambda$ as shown in Supp. Fig.1(a)(iii), and corresponds to the maximum similarity distance between the reconstructed and true densities. We show the reconstructed densities at these three strengths in Supp. Fig.1(a)(iv-vi), and remark that $\lambda=1.22$ shows noise artifacts inside and outside the wire and is thus before the turn in the $l-$curve. The $\lambda=2.04$ shows no artifacts, but has a higher offset. Instead, the maxima of the SSIM curve at $\lambda=2.53$ shows a lower offset with no artifacts outside the wire. This demonstrates that the $l-$curve is not exact for TV regularization and should instead be used in supplement with SSIM to choose the regularizer strength.
@@
The outcomes for a high standoff are presented in Supp. Fig.1(b). The $l$-curve exhibits a distinct inflection point at $\lambda=0.2$, with two supplementary points $\lambda\in\{0.06,0.41\}$ marked for reference. However, the SSIM curve displays a noticeable contrast compared to the low standoff scenario. Instead of a peak, there is a plateau, and a reference local maximum is identified on this plateau at $\lambda=0.41$. Additionally, as anticipated, the SSIM distance is lower than in the low standoff case. This indicates that the SSIM curve alone is insufficient for regularizer selection, as all models with $\lambda\in[0.2,0.5]$ yield the same SSIM, rendering it challenging to make a selection. This issue is further elucidated by the reconstructions in Supp. Fig.1(b)(iv-vi). The image corresponding to $\lambda=0.2$, identified by the kink in the $l$-curve, demonstrates the optimal balance between resolving wire features and suppressing noise. In contrast, the $\lambda=0.06$ image exhibits excessive noise, while the $\lambda=0.41$ plot shows significant broadening. Consequently, employing the $l$-curve in conjunction with an additional distance metric like SSIM offers a reliable method for selecting the regularizer strength in Bayesian optimization.